# MA4DIV: Multi-Agent Reinforcement Learning for Search Result Diversification

### v2

@@ 5.2. Experimental Settings
We compare MA4DIV with several state-of-the-art baselines in search result diversification, including the heuristic methods:
5.3. Experimental Results
Table 3 and 4 report the performances of all baselines and our MA4DIV in terms of the six diversity performance metrics, including $\alpha$-NDCG@5, $\alpha$-NDCG@10, ERR-IA@5, ERR-IA@10, S-recall@5, S-recall@10 on the TREC web track datasets and the DU-DIV dataset respectively. We highlight the top three algorithms on each metric by darkening the background color of the numbers.
5.3.1. TREC web track datasets
@@
As mentioned in Section 5.1, TREC web track datasets have only 198 unique queries, deep learning methods are prone to overfitting on this dataset. Figure 2 shows the training curve of MA4DIV on the TREC web track datasets. We can see, after about the 11 iterations, MA4DIV achieves the best performance on the test data, and the subsequent iterations fall into overfitting the relatively small training set. However, the evaluation metric $\alpha$-NDCG@10 on the training set grows to more than 0.8. These results indicate that:
  • Our MA4DIV is capable to effectively optimize the diversity metric $\alpha$-NDCG@10.
  • However, it could easily overfit the relatively small training set of the TREC web track dataset, which only contains 198 unique queries.
  • Therefore, MA4DIV has the potential to perform even better in the search result diversification task when given a larger dataset.
Therefore, we further construct the DU-DIV dataset and conduct extensive experiments on it.
@@
5.3.2. DU-DIV dataset
Table 4 shows results in different evaluation metrics on the DU-DIV dataset. Obviously, MA4DIV achieves state-of-the-art performances on all six evaluation metrics. Heuristic algorithms, MMR and xQuAD, perform poorly on $\alpha$-NDCG. Comparing to single-agent reinforcement learning methods, MDP-DIV and $\mathrm{M^{2}Div}$, MA4DIV has a better performance on evaluation metrics which demonstrate that MA4DIV can improve the diversity of search results. And MA4DIV also outperforms DALETOR and MO4SRD which optimize the approximate value of the evaluation metrics. This verifies that directly optimizing the evaluation metrics is indeed better than optimizing the approximation of the evaluation metrics. All these results demonstrate that the proposed MA4DIV method outperforms existing search result diversification methods on an industrial scale dataset (answering RQ.2).
Interestingly, we notice that MDP-DIV, which does not perform well on the TREC web track datasets, is the second best performing algorithm on the DU-DIV dataset, following our MA4DIV in terms of $\alpha$-NDCG@5 and $\alpha$-NDCG@10. In MDP-DIV, the reward function is defined as: $R(s_{t},a_{t})=\alpha\textrm{-}DCG[t+1]-\alpha\textrm{-}DCG[t]$, where $\alpha\textrm{-}DCG[t]$ is the discounted cumulative gain at the $t$-th ranking position. And the objective function of MDP-DIV is to maximize the cumulative rewards in one episode, that is $G_{t}=\sum_{k=0}^{n-1-t}\gamma^{k}R_{t+k+1}$. Consider that Xia et al. (2017) sets $\gamma\textrm{=}1$ and $\alpha\textrm{-}DCG[t]\textrm{=}0$, the objective function can be rewritten as $G_{t}=\alpha\textrm{-}DCG[n]$, which is equal to $\alpha\textrm{-}DCG$ defined in Equation (2). In addition, the objective function of our MA4DIV defined in Equation (10) is equal to $\alpha\textrm{-}NDCG$. Essentially, the optimization objective of MDP-DIV is basically the same as that of our MA4DIV. Therefore, due to the correctness of its optimization objective, MDP-DIV can outperform all the other baselines.
@@ 5.4. Comparison and Analysis of Efficiency
5.4.1. Training Time on TREC web track datasets
To highlight the efficiency of our MA4DIV, we also compare several algorithms from the perspective of training time. Table 5 shows the comparison of training time taken by different algorithms to achieve their optimal performance on the TREC web track datasets. To ensure a fair comparison, all the algorithms in Table 5 run on the same machine. We can see that $\mathrm{M^{2}Div}$ takes over a day to reach optimal performance. This is because $\mathrm{M^{2}Div}$ utilizes Monte Carlo Tree Search for policy enhancement, which is highly time-consuming. The time taken by MDP-DIV and MO4SRD is close and significantly less than $\mathrm{M^{2}Div}$. The proposed MA4DIV takes the shortest time, approximately only 20$\sim$25 minutes. This is due to the efficient exploration ability of using a multi-agent to model the ranking process, in which a permutation of the given documents set ($n$ documents) can be obtained at one time step. The single agent reinforcement learning algorithm MDP-DIV can only explore a permutation of the given documents after a whole episode which contains $n$ time steps. This is the main reason why the training time of MA4DIV is significantly shorter than MDP-DIV. In MO4SRD, the complexity of calculating all approximate ranking positions $r_{i}$ in Equation (1) is $\mathcal{O}(n^{2})$ and the complexity of calculating all $c_{li}$ approximations is $\mathcal{O}(n^{2}m)$, which are time-consuming parts. While MA4DIV directly optimizes Equation (2) without complex approximations, which takes significantly less time.
@@ 5.4.2. Iteration Times on DU-DIV dataset
As the optimization goals of MA4DIV and MDP-DIV are essentially the same as we discussed in Section 5.3.2, we are interested in comparing the learning efficiency of the proposed multi-agent RL approach and the existing single-agent RL method.
@@
The main reason for the results just mentioned is as follows: MDP-DIV adopts an on-policy RL method to optimize the cumulative rewards. A feature of on-policy RL is that the currently sampled data is discarded after it is utilized. While MA4DIV is a off-policy RL method. From 9-th line and 12-th line of Algorithm 2, we can see that the sampled data is stored in replay buffer $\mathcal{M}$, then update the model multiple times in minibatch approach. In addition, each data stored in replay buffer $\mathcal{M}$ may be sampled multiple times to update the ranking model. Therefore, the data utilization efficiency of MA4DIV is higher than that of MDP-DIV, which is the main reason why MA4DIV converges significantly faster.
The shorter training time on the TREC web track datasets and the faster convergence rate on the DU-DIV dataset jointly demonstrate that MA4DIV has high efficiency in training, which answers RQ.3.
Figure 3. Training curves of MA4DIV and MDP-DIV on DU-DIV dataset.
@@ 6. Conclusions and Future Work
In this paper we have proposed a new method for search result diversification, referred to as MA4DIV. To alleviate the sub-optimal problem and improve the efficiency of training, we consider each document as an agent and model the diversity ranking process as a multi-agent cooperative task. We conduct experiments on the TREC web track dataset to prove the effectiveness and potential of our MA4DIV. Furthermore, we construct a large-scale dataset called DU-DIV using the real data from Baidu search engine. And MA4DIV achieve state-of-the-art performance and high learning efficiency on the DU-DIV dataset.

### v3

@@ 5.2. Experimental Results
5.2.1. The Results on TREC web track datasets
As mentioned in Table 1, TREC web track datasets have only 198 unique queries, deep learning methods are prone to overfitting on this dataset. Figure 2 shows the training curve of MA4DIV on the TREC web track datasets. We can see, after about the 11 iterations, MA4DIV achieves the best performance on the test data, and the subsequent iterations fall into overfitting the relatively small training set. However, the evaluation metric $\alpha$-NDCG@10 on the training set grows to more than 0.8. These results indicate that: (1) Our MA4DIV is capable to effectively optimize the diversity metric $\alpha$-NDCG@10. (2) It could easily overfit the relatively small training set of the TREC web track dataset, which only contains 198 unique queries. (3) MA4DIV has the potential to perform better in the search result diversification task when given a larger dataset.
@@ 5.2.1. The Results on TREC web track datasets
Therefore, we construct a larger diversity dataset called DU-DIV from a real industrial search engine. And then we show the experiments on DU-DIV dataset.
5.2.2. The Results on DU-DIV dataset
Table 3 shows results in different evaluation metrics on the DU-DIV dataset. MA4DIV achieves state-of-the-art performances on all six evaluation metrics. Heuristic algorithms, MMR and xQuAD, perform poorly on $\alpha$-NDCG. Comparing to single-agent reinforcement learning methods, MDP-DIV and $\mathrm{M^{2}Div}$, MA4DIV has a better performance on evaluation metrics which demonstrate that MA4DIV can improve the diversity of search results. MA4DIV also outperforms DALETOR and MO4SRD which optimize the approximate value of the evaluation metrics. This verifies that directly optimizing the evaluation metrics is indeed better than optimizing the approximation of the evaluation metrics. All these results demonstrate that the proposed MA4DIV method outperforms existing search result diversification methods on an industrial scale dataset (answering RQ.2).
Interestingly, we notice that MDP-DIV, which does not perform well on the TREC web track datasets, is the second best performing algorithm on the DU-DIV dataset, following our MA4DIV in terms of $\alpha$-NDCG@5 and $\alpha$-NDCG@10. In MDP-DIV, the reward function is defined as: $R(s_{t},a_{t})=\alpha\textrm{-}DCG[t+1]-\alpha\textrm{-}DCG[t]$, where $\alpha\textrm{-}DCG[t]$ is the discounted cumulative gain at the $t$-th ranking position. And the objective function of MDP-DIV is to maximize the cumulative rewards in one episode, that is $G_{t}=\sum_{k=0}^{n-1-t}\gamma^{k}R_{t+k+1}$. Consider that Xia et al. (Xia et al., 2017) sets $\gamma\textrm{=}1$ and $\alpha\textrm{-}DCG[t]\textrm{=}0$, the objective function can be rewritten as $G_{t}=\alpha\textrm{-}DCG[n]$, which is different from $\alpha\textrm{-}NDCG$ with just a normalization factor. In addition, the objective function of our MA4DIV defined in Equation (5) is equal to $\alpha\textrm{-}NDCG$. Similar to MA4DIV, MDP-DIV also directly optimizes one of the end evaluation metrics in SRD. Therefore, this provides a theoretical analysis for MDP-DIV to be able to outperform other baselines.
@@ 5.3. Comparison and Analysis of Efficiency
5.3.1. Training Time on TREC web track datasets
To highlight the efficiency of our MA4DIV, we also compare several algorithms from the perspective of training time. Table 4 shows the comparison of training time taken by different algorithms to achieve their optimal performance on the TREC web track datasets. To ensure a fair comparison, all the algorithms in Table 4 run on the same machine.
@@ 5.3.1. Training Time on TREC web track datasets
We can see that $\mathrm{M^{2}Div}$ takes over a day to reach optimal performance. This is because $\mathrm{M^{2}Div}$ utilizes Monte Carlo Tree Search for policy enhancement, which is highly time-consuming. The time taken by MDP-DIV and MO4SRD is close and significantly less than $\mathrm{M^{2}Div}$. MA4DIV takes the shortest time, only 20$\sim$25 minutes. This is due to the efficient exploration ability of using a multi-agent to model the ranking process, in which a permutation of the given documents set ($n$ documents) can be obtained at one time step. The single agent reinforcement learning algorithm MDP-DIV can only explore a permutation of the given documents after a whole episode which contains $n$ time steps. This is the main reason why the training time of MA4DIV is significantly shorter than MDP-DIV. In MO4SRD, the complexity of calculating all approximate ranking positions $r_{i}$ in $\alpha\textrm{-}DCG$ is $\mathcal{O}(n^{2})$ and the complexity of calculating all $c_{li}$ approximations is $\mathcal{O}(n^{2}m)$, which are time-consuming parts. While MA4DIV directly optimizes $\alpha\textrm{-}NDCG$ without complex approximations, which takes significantly less time.
@@ 5.3.2. Iteration Times on DU-DIV dataset
As the optimization goals of MA4DIV and MDP-DIV are essentially the same (we discussed in the analysis of the results on DU-DIV dataset), we are interested in comparing the learning efficiency of the proposed multi-agent RL approach and the existing single-agent RL method.
Figure 3. Training curves of MA4DIV and MDP-DIV on DU-DIV dataset.
@@
The main reason for the results just mentioned can be seen in Appendix F.
5.3.3. Analysis of Inference Complexity
The analysis of the inference complexity about different diversified search algorithms can be seen in Appendix G.
The Section 5.3 shows that MA4DIV have the shorter training time on the TREC web track datasets, the faster convergence rate on the DU-DIV dataset and the lowest inference complexity. And these advantages jointly demonstrate that MA4DIV has high efficiency in training and inference process, which answers RQ.3.
@@ 6. Conclusions and Future Work
In this paper we have proposed a novel method for search result diversification using MARL, called MA4DIV. To alleviate the sub-optimal problem and improve the efficiency of training, we consider each document as an agent and model the diversity ranking process as a multi-agent cooperative task. We conduct experiments on the TREC web track dataset to prove the effectiveness and potential of our MA4DIV. Furthermore, we construct a larger scale dataset called DU-DIV using the real data from industrial search engine. And MA4DIV achieve state-of-the-art performance and high learning efficiency on the DU-DIV dataset.
@@
Acknowledgments
This research was supported by the Natural Science Foundation of China (61902209, 62377044, U2001212), and Beijing Outstanding Young Scientist Program (NO.BJJWZYJH012019100020098), Intelligent Social Governance Platform, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Renmin University of China, Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education, and the fund for building world-class universities (disciplines) of Renmin University of China.
Appendix
Appendix A Why do existing methods lead to suboptimal diversity ranking?
A.1. Why the ”greedy selection” approach leads to suboptimality.
In the task of Search Result Diversification, the “greedy selection” methods (Carbonell and Goldstein, 1998; Santos et al., 2010; Xia et al., 2017; Feng et al., 2018) typically refer to selecting the document that can provide the most additional information at each step when generating a diversified result list, i.e., making a locally optimal choice. However, this method is prone to suboptimal rankings for the following reasons:
  • Local Optimality Instead of Global Optimality: The greedy selection method chooses the document most beneficial to the current ranking at each step, but this method only considers the optimal solution for the current step, not the global optimality of the entire ranking list. Hence, although each step is a locally optimal choice, the final result may not be the best overall ranking.
  • Lack of Holistic Perspective: As the greedy selection is performed step by step, it does not take into account all possible combinations of documents and their overall effects. This might result in missing out on document combinations that could provide broader topic coverage.
  • Ignoring Interactions Among Documents: Greedy selection often does not fully consider the interrelationships among documents. For instance, some documents might be highly related in content, and their consecutive appearance might reduce user satisfaction and the diversity of the results.
  • Mismatch Between Objective Function and Evaluation Metrics: The objective function used in the training of the greedy selection method may not match the metrics used in the final evaluation. This inconsistency could lead to optimising for a target in the training process that does not match the performance needing evaluation.
  • Lack of Flexibility: Greedy selection methods are usually not flexible; once the ranking order is determined, it is difficult to adjust. If the selected documents are not optimal, i.e., there is an error in the ranking compared to the optimal one, then the documents chosen based on the currently selected documents will result in cumulative errors.
A.2. Why optimizing an approximation of the objective function leads to suboptimality.
Methods of optimizing the approximation of the objective function (Yan et al., 2021; Yu, 2022) can often achieve parallel inference (faster than the serial selection of “greedy selection”), but may result in suboptimal ranking results due to the following issues:
  • Mismatch Between Objective Function and Evaluation Metrics: There is a mismatch between the optimization objective used in the training process and the evaluation metrics used in the final assessment, which may prevent the model from accurately learning to achieve the best diversification effect.
  • Low Exploration Efficiency: Such methods are less capable of exploring different document ranking combinations. When faced with the challenge of exploring as many different document ranking combinations as possible in a vast combination space, it leads to an increase in the time complexity of training and makes it difficult to find the global optimum.
Appendix B Diversity Evaluation Metrics
B.1. $\alpha\textrm{-}NDCG$
B.2. $ERR\textrm{-}IA$
The $ERR\textrm{-}IA$ is defined in Equation (18).
B.3. $S\textrm{-}recall$
Appendix C The Detail of MHSA
C.1. The structure of MHSA
MHSA further enhances the model capacity by applying the above process multiple times in parallel, which is called multi-head mechanism. Suppose there are $h$ heads, the output after multi-head is $\mathbf{O}=[\mathbf{O}_{1};\mathbf{O}_{2};\cdots;\mathbf{O}_{h}]\mathbf{W}_{O}$, where []delimited-[][$\cdot$][ ⋅ ] denotes concatenation, and $\mathbf{O}_{i}$ and $\mathbf{W}_{O}$ are the output of the $i$-th head and a learnable weight matrix with dimensions $h\cdot d_{v}\times d$, respectively.
C.2. The reason for using MHSA
According to (Pang et al., 2020), the ranking model should satisfy the permutation invariance property which is described as following Definition:
Definition 0 (Permutation Invariance).
A method to construct a ranking model in accordance with Permutation Invariance involves initially attributing scores to the documents via a permutation equivariant scoring function, followed by ordering based on these scores. Pang et al. (Pang et al., 2020) also proved that the multi-head attention block is permutation equivariant. Therefore, the MHSA module is used to compute higher-level cross feature vectors between documents.
Appendix D The Introduction of Datasets
We conduct experiments to address above research questions on two datasets, TREC 2009$\sim$2012 Web Track datasets and DU-DIV dataset. TREC 2009$\sim$2012 Web Track datasets are publicly available and Xia et al. (Xia et al., 2017) first uses them for training and evaluating diversified search models. Almost all the following works, such as $\mathrm{M^{2}Div}$ (Feng et al., 2018), DALETOR (Yan et al., 2021), MO4SRD (Yu, 2022), etc., conduct experiments on these datasets. As a widely used test collection for diversified search, the TREC Web track dataset only contains 198 queries, and previous studies often use Doc2vec (Le and Mikolov, 2014) to obtain the document embeddings on this dataset. In this work, to ensure a fair and meaningful comparison with existing work, we first compare MA4DIV with existing baselines on the this relatively small dataset. Then, we further conduct extensive experiments on a new and larger scale diversity dataset called DU-DIV, which is constructed based on real user data in industrial search engine. The Figure 4 shows the average number of subtopics of the 15 documents with ideal ranking permutation in the DU-DIV dataset.
Figure 4. Average number of subtopics across the 15 documents in the DU-DIV dataset. The abscissa from 1 to 15 represents positions 1 to 15 in the ideal permutation, and the ordinate is the average of the number of subtopics contained in the documents of the corresponding position.
Table 5. Comparisons of datasets TREC and DU-DIV.
Dataset TREC DU-DIV
Source TREC 2009$\sim$2012 Web Track Real search engine
No. of Queries 198 4473
No. of Documents per Query average 211 docs/query 15 docs/query
No. of Subtopics maximum 7 subtopics/doc total 50 subtopics
Representation Doc2vec (Le and Mikolov, 2014) BERT (Devlin et al., 2018)
More details of these two datasets are shown in Table 5. The DU-DIV dataset has 4473 queries, far more than the TREC datasets. The documents in the DU-DIV dataset are the top-15 documents retrieved by the industrial search engine for each query, which makes this dataset particularly useful for the diversified ranking task in which we need to rerank these top documents for diversity. In the TREC web track datasets, each document contains a maximum of 7 subtopics. In the DU-DIV dataset, we predefined 50 subtopics, such as law, healthcare, education, etc. Each document corresponds to one or more of the predefined subtopics. Moreover, while previous studies on the TREC web track dataset often use the Doc2vec model to represent queries and documents in vector with $L\textrm{=}100$ dimensions, the DU-DIV dataset utilizes the more powerful model BERT (Devlin et al., 2018) to encode queries and documents into vectors with $L\textrm{=}1024$ dimensions.
Appendix E Experimental Settings
Considering that TREC datasets contain only 198 queries, this is prone to overfitting for existing deep learning models. In addition, the average number of documents related to each query is about 211, but a large number of these documents do not contain any subtopics. Therefore, we reshape TREC dataset. That is, from each of the original document lists in TREC (198 lists in total, with an average of 211 documents per query), we subsample 30 documents at a time to form a new document list. This process runs multiple times over a document list associated with the same query, resulting in a reshaped dataset with a total of 6,232 document lists (198 unique queries). And it’s crucial to ensure that an appropriate number of the 30 documents sampled contain at least one subtopic. Figure 5 shows the number of documents with subtopics in these 6,232 document lists. We can see that there are 1,053, 1,717, and 1,980 document lists with 16 $\sim$ 20, 21 $\sim$ 25, 26 $\sim$ 30 documents containing any subtopics, and 758 document lists with 1 $\sim$ 5 documents containing any subtopics. This means that after our sampling, most document lists contain a rich number of subtopics, and some document lists have a small number of subtopics. Overall, this ensures that the data is comprehensive.
Figure 5. This figure illustrates the relationship between the quantity of newly sampled document lists and the count of documents that include any subtopics present in the sampled list.
Table 6. Hyperparemeters of MA4DIV on different dataset.
Dataset $|\mathcal{A}|$ $L$ $H$ $z$
TREC 30 1 4 64
DU-DIV 15 1 4 256
Appendix F Reason for Higher Efficiency of MA4DIV
The main reason for the higher efficiency of MA4DIV is as follows: MDP-DIV adopts an on-policy RL method to optimize the cumulative rewards. A feature of on-policy RL is that the currently sampled data is discarded after it is utilized. While MA4DIV is a off-policy RL method. From 9-th line and 12-th line of Algorithm 2, we can see that the sampled data is stored in replay buffer $\mathcal{M}$, then update the model multiple times in minibatch approach. In addition, each data stored in replay buffer $\mathcal{M}$ may be sampled multiple times to update the ranking model. Therefore, the data utilization efficiency of MA4DIV is higher than that of MDP-DIV, which is the main reason why MA4DIV converges significantly faster.
Appendix G Analysis of Inference Complexity
Table 7 shows the inference complexity of different approaches. We can see that the Heuristic Approaches, including MMR (Carbonell and Goldstein, 1998) and xQuAD (Santos et al., 2010), have to perform inference process in $O(n)$. And $n$ is the number of candidate documents. In addition, Single-Agent RL Approaches, such as MDP-DIV (Xia et al., 2017), also has a $O(n)$ inference complexity. These two approaches can only select one document in one time step, so the inference complexity is equal to the number of candidate documents.
Because DALETOR (Yan et al., 2021) and MO4SRD (Yu, 2022) optimize the approximation of evaluation metrics and can score for all candidate documents at once, these approaches have $O(1)$ inference complexity. In our MA4DIV, the ranking scores $\{s_{1},\dots,s_{n}\}$ can be obtained in one time step, so MA4DIV also has the lowest inference complexity with $O(1)$.
Table 7. The inference complexity of different diversified search algorithms. $n$ is the number of candidate documents.
Method Inference Complexity
Heuristic Approaches $O(n)$
Single-Agent RL Approaches $O(n)$
Approximate Metric Approaches $O(1)$
MA4DIV (ours) $O(1)$